OpenStudy (anonymous):

what is the factored form of the expression? w^2 +12w + 36

OpenStudy (anonymous):

(w+6)(w+6) or (w+6)^2

OpenStudy (anonymous):

\[w^2 +12w + 36 =w^2 +6w +6w+ 36=w(w +6)+6(w+ 6)=(w+6)(w+6)\] @ThatBoyIsANinja